# Local map Construction Methods with SD map: A Novel Survey

### v1

@@ Abstract
In recent years, significant academic advancements have been made in the field of autonomous vehicles, with Local maps emerging as a crucial component of autonomous driving technology. Local maps not only provide intricate details of road networks but also serve as fundamental inputs for critical tasks such as vehicle localization, navigation, and decision-making. Given the characteristics of SD map (Standard Definition Map), which include low cost, ease of acquisition, and high versatility, perception methods that integrate SD map as prior information have demonstrated significant potential in the field of Local map perception.
The purpose of this paper is to provide researchers with a comprehensive overview and summary of the latest advancements in the integration of SD map as prior information for Local map perception methods. This review begins by introducing the task definition and general pipeline of local map perception methods that incorporate SD maps as prior information, along with relevant public datasets. And then it focuses on the representation and encoding methods of multi-source information, as well as the methods for fusing multi-source information. In response to this burgeoning trend, this article presents a comprehensive and meticulous overview of the diverse research efforts in this particular field. Finally, the article addresses pertinent issues and future challenges with the aim of guiding researchers in understanding the current trends and methodologies prevalent in the field.
@@ 1 Introduction
Local map perception is a critical and challenging task in the field of intelligent driving. It involves detailed understanding and real-time modeling of the environment surrounding the vehicle, serving as the foundation for decision-making and navigation in autonomous systems. Local maps not only provide information about roads and lanes but also encompass the detection and recognition of obstacles, traffic signs, pedestrians, and other dynamic or static objects. This information is essential for ensuring the safe operation of the vehicle and efficient path planning. Without precise local map perception, autonomous vehicles may deviate from their routes, cause traffic accidents, and potentially endanger passenger safety. Thus, Local map perception plays an indispensable role in the autonomous driving ecosystem.
Unlike typical object detection, Local map perception must handle complex and dynamic environmental information while maintaining high accuracy under various lighting conditions and weather situations. For example, shadows on the road, light reflections, dynamic obstacles, and occlusions of road signs can all interfere with Local map perception.
Additionally, sensor noise and data delays further complicate the perception task. Therefore, developing robust Local map perception technologies is crucial for achieving safe and reliable autonomous driving. To solve this kind of problem, many researchers have proposed various methods. Chen and Lei [1] proposed a method for visual localization and map construction using ground texture, enhancing the accuracy of positioning and the precision of map updates through global and local optimization. [2] enhanced online map prediction and lane topology understanding by leveraging SD maps and integrating SD map information through a Transformer encoder, which alleviates problems with obscured lane lines or poor visibility and significantly improves the performance of lane detection and topology prediction. [3] proposed an innovative video lane detection algorithm that enhances the feature map of the current frame using an occlusion-aware memory-based refinement (OMR) module, leveraging obstacle masks and memory information to improve the detection accuracy and robustness under occlusion. RVLD[4] improved the reliability of lane detection by recursively propagating the state of the current frame to the next frame, utilizing the information from the previous frames. Besides, there are Laneaf[5], LaneATT[6], Streammapnet[7] to ease these issues.
In previous autonomous driving research, high-definition maps (HDMap) have been vital. HDMap, characterized by absolute and relative accuracies within 1 meter, provides high precision, freshness, and richness of electronic maps, including extensive road and environmental information. These maps offer precise navigation and positioning services to support safe and efficient autonomous driving. However, HDMap faces significant challenges, primarily in terms of real-time updates and cost control. Urban road environments change frequently, and any minor alteration can impact the driving safety of autonomous vehicles. Traditional HDMap production methods require substantial time and resources, making real-time updates difficult, both [8] and [9] have pointed out similar issues. Moreover, the production and maintenance costs of HDMap are extremely high, with costs reaching thousands of dollars per kilometer using traditional methods.
@@
In this context, the ”highly rely on perception, reducing dependence on maps” approach has gained widespread recognition within the industry. This approach emphasizes the use of onboard sensors for autonomous driving perception tasks, supplemented by lightweight map information. This strategy reduces reliance on real-time map updates, lowering maintenance costs, while lightweight map information can effectively compensate for certain limitations of onboard sensors, enhancing the model’s robustness. SD map, as a widely used electronic map in traffic navigation and geographic information services, feature low production and maintenance costs, easy accessibility, and small data size, making it suitable as lightweight maps to assist onboard sensors in constructing local maps for autonomous driving.
@@
Despite the promising prospects and numerous challenges in constructing Local maps based on SD map, there is a lack of comprehensive research reviews in this area. To address this gap, this review aims to provide a thorough overview of the latest advancements in Local map construction methods utilizing SD map. Specifically, the focus is on the application of SD map information representation methods and multimodal data fusion techniques in local map perception tasks. This research delves into the major developments, challenges, and research directions in this field. The contributions to the existing body of knowledge are as follows, The existing literature on local map construction using SD map as a prior is reviewed. The strengths and limitations of these approaches are analyzed, providing insights into their effectiveness and applicability in real-time autonomous driving applications. The representation and encoding methods of information from various sensors, as well as the fusion techniques for multi-source sensor data, are highlighted, which are crucial for real-time local map generation. The underlying principles, architecture, and performance of these methods are discussed, shedding light on their feasibility and practicality in the field key challenges and open research questions in local map construction using SD map as a prior are identified.
@@
Figure 1: General pipeline of local map construction with SD map
2 Background
@@
2.1 Task Definition of Local map Construction with SD Map
The task of Local map perception involves creating an accurate map representing the vehicle’s surrounding environment to support autonomous driving decision-making and planning. This task typically relies on data from various sensors, including cameras, LiDAR, radar, and GPS. Additionally, incorporating prior information from SD map enhances the model’s robustness and mitigates the impact of uncertainties from onboard sensors, improving the overall model performance. The core of the Local map perception task is real-time sensing and understanding of the vehicle’s surroundings.
@@ 2.2 Standard Definition Map
SD map, short for Standard Definition Map, is a digital map technology providing basic geographic information and road network structures. It is widely used in everyday navigation and geographic information services, offering convenience to users. SD map primarily provides the centerline skeleton of roads without detailed lane information, road signs, or other high-precision environmental features. For the task of Local map construction, SD map offers three main advantages. First, SD map data is easily accessible. It can typically be obtained for free from open geographic data sources such as OpenStreetMap [10], making it suitable for large-scale applications. Second, compared to HDMap, the production and maintenance costs of SD Map are significantly lower. Lastly, SD map has high universality, covering most types of roads, and can provide relevant road information for local map construction tasks. Platforms like OSM and Baidu Maps can serve as data sources for SD map. For instance, OpenStreetMap (OSM) is a collaborative project created and maintained by global volunteers, providing free, editable, and open-content maps. OSM data includes a wide range of geographic information such as roads, buildings, parks, and rivers, which users can freely access, edit, and use.
@@ 2.3 Datasets
In the field of BEV Local map construction, commonly used datasets include KITTI[11], nuScenes[12], ApolloScape[13], Argoverse[14], Openlane[15], and Waymo[16] Open Dataset.
@@
The KITTI[11] dataset, created by the Karlsruhe Institute of Technology and the Toyota Technological Institute, provides stereo camera, LiDAR, and GPS/IMU data, covering urban, rural, and highway scenes, and is suitable for tasks such as object detection, tracking, and road detection.
nuScenes[12], released by Motional, includes data from six cameras, five radars, one LiDAR, IMU, and GPS, suitable for urban traffic scenarios under various weather and lighting conditions.
ApolloScape[13], released by Baidu, offers high-precision 3D annotated data covering various urban road scenes, suitable for tasks like lane detection and semantic segmentation.
@@
Argoverse[14], released by Argo AI, includes stereo camera, LiDAR, GPS, and IMU data, providing detailed 3D annotations and lane markings, primarily used for 3D object detection and lane detection.
@@
The Waymo[16] Open Dataset, released by Waymo, covers a variety of weather and traffic conditions, providing high-quality data from LiDAR and cameras, suitable for tasks such as 3D object detection, tracking, and lane detection.
@@
OpenLane-V2[17], also known as OpenLane-Huawei or Road Genome, is a benchmark dataset for next-generation autonomous driving scene road structure perception, jointly open-sourced by Shanghai Artificial Intelligence Laboratory and Huawei Noah’s Ark Laboratory. It is the first dataset to include the topological relationships of road structures in traffic scenes.
@@
ONCE-3DLanes[18] dataset, a real-world autonomous driving dataset with lane layout annotation in 3D space, is a new benchmark constructed to stimulate the development of monocular 3D lane detection methods. It is collected in various geographical locations in China, including highway, bridge, tunnel, suburb and downtown, with different weather conditions (sunny/rainy) and lighting conditions (day/night). The whole dataset contains 211K images with its corresponding 3D lane annotations in the camera coordinates.
@@
CurveLanes[19] is a new benchmark lane detection dataset with 150K lanes images for difficult scenarios such as curves and multi-lanes in traffic lane detection. It is collected in real urban and highway scenarios in multiple cities in China. All images are carefully selected so that most of them image contains at least one curve lane. More difficult scenarios such as S-curves, Y-lanes, night and multi-lanes can be found in this dataset.
@@ 2.4.1 Metrics for Lane Extraction
In this formula, mAP represents the mean Average Precision, measuring detection accuracy. The TP set contains the average values of five True Positive Metrics: ATE (Average Translation Error), ASE (Average Scale Error), AOE (Average Orientation Error), AVE (Average Velocity Error), and AAE (Average Attribute Error).
@@ 2.4.2 Metrics for Topology Reasoning
OpenLane-V2[17] breaks down the task into three subtasks: 3D lane detection, traffic element recognition, and topology reasoning. The overall task performance is described using the OpenLane-V2 Score (OLS), which is the average of the metrics for each subtask. The metric for 3D lane detection, DETl, can be expressed as the mean AP at different thresholds $t\in T$, $T$={1.0,2.0,3.0} [Equation], where AP is calculated using the Fréchet distance. Traffic element detection is evaluated similarly to object detection using AP, with an IoU threshold set to 0.75. Traffic elements have various attributes, such as traffic light colors, which are closely related to lane accessibility, so attributes must also be considered. Assuming A is the set of all attributes, the evaluation includes attribute classification accuracy in.
$\text{OLS}=\frac{1}{4}\left[\text{DET}_{l}+\text{DET}_{t}+\text{f}(\text{TOP}_{ll})+\text{f}(\text{TOP}_{lt})\right]$ (4)
@@
OpenLane-V2 uses the TOP score to evaluate the quality of topology reasoning, akin to the mAP metric but adapted for graphs. Essentially, this converts the topology prediction problem into a link prediction problem and calculates mAP (mean APs of all vertices) to assess algorithm performance. The first step is to determine a matching method to pair ground truth and predicted vertices (i.e., centerlines and traffic elements). For centerlines, Fréchet distance is used; for traffic elements, IoU is used. When the confidence score of an edge between two vertices exceeds 0.5, they are considered connected. The vertex AP is obtained by ranking all predicted edges of a vertex and calculating the mean of cumulative precision:

### v2

# Local Map Construction with SDMap: A Comprehensive Survey

@@ Abstract
Local map construction is a vital component of intelligent driving perception, offering necessary reference for vehicle positioning and planning. Standard Definition map (SDMap), known for its low cost, accessibility, and versatility, has significant potential as prior information for local map perception. This paper mainly reviews the local map construction methods with SDMap, including definitions, general processing flow, and datasets. Besides, this paper analyzes multimodal data representation and fusion methods in SDMap-based local map construction. This paper also discusses key challenges and future directions, such as optimizing SDMap processing, enhancing spatial alignment with real-time data, and incorporating richer environmental information. At last, the review looks forward to future research focusing on enhancing road topology inference and multimodal data fusion to improve the robustness and scalability of local map perception.
@@ 1 Introduction
In intelligent driving, the construction of a local map is crucial for providing real-time awareness of the vehicle’s surroundings, enabling precise navigation. It reflects the current state of the environment using sensor data and helps in predicting the movement of surrounding objects. By offering real-time updates, local maps support dynamic decision-making and control, even in unmapped areas. They enhance robustness, reduce computational load, and ensure that the vehicle can safely adapt to rapidly changing conditions. Local maps provide accurate static information such as road geometries and traffic sign locations, serving as prior knowledge that enhances the robustness of perception systems in complex environments. Local map perception systems are expected to achieve high-precision detection in all weather and lighting conditions, as any instance of missed or incorrect detection can result in significant consequences.
Figure 1: The overview of local map construction with SDMap.
@@
High-definition map (HDMap) is a vital topic in previous research. HDMap provides high precision, freshness, and richness of electronic map. However, HDMap faces significant challenges, primarily in terms of real-time updates and cost control. Urban road environments change frequently, and any minor alteration can impact the driving safety of autonomous vehicles. Traditional HDMap production methods require substantial time and resources, making real-time updates difficult. By heavily relying on perception and reducing dependence on detailed maps, this approach has gained widespread recognition. It emphasizes the use of onboard sensors for intelligent driving perception tasks, supplemented by lightweight map information. This strategy reduces reliance on real-time map updates, lowering maintenance costs, while lightweight map information can effectively compensate for certain limitations of onboard sensors, enhancing the model’s robustness. SDMap, as a widely used electronic map in traffic navigation and geographic information services, features low production and maintenance costs, easy accessibility, and small data size, making it suitable as lightweight map to assist onboard sensors in constructing local map for intelligent driving.
@@
Despite the promising prospects and numerous challenges in constructing local map based on SDMap, there is a lack of comprehensive research reviews in this area. To address this gap, this review aims to provide a thorough overview of the latest advancements in local map construction methods utilizing SDMap. Specifically, the focus is on the application of SDMap information representation methods and multimodal data fusion techniques in local map construction tasks. This research delves into the major developments, challenges, and research directions in this field. The contributions to the existing body of knowledge are as follows, The existing literature on local map construction using SDMap as a prior is reviewed. The representation and encoding methods of information from various sensors, as well as the fusion techniques for multimodal data. The underlying principles, architecture, and performance of these methods are discussed, shedding light on their feasibility and practicality in the field key challenges and open research questions in local map construction using SDMap as a prior are identified.
Fig. 1 presents the framework and context of the comprehensive review, where the relationships between the modules in the figure are organized in order and hierarchy. This paper analyzes and organizes from the perspective of processing multimodal data. The data is categorized into images, lidar point clouds, SDMap data, and multimodal data. Each category corresponds to specific processing and fusion methods, with distinct techniques and formats for handling and integration. It can be seen in the figure that in the local map construction method with SDMap, the image processing methods are divided into CNN-based modules or transformer-based modules, while the point cloud data is divided into extraction processing methods and fusion methods with image data. The processing methods for SDMap data are divided into Raster and Vector categories based on the form of processed data. Finally, the processing methods for multimodal data are divided into spatial alignment and feature fusion, the former is further divided into SLAM-based and attention-based methods, while the latter fusion methods are divided into addition, concatenation, and cross-attention methods. This classification structure helps to clearly sort out the different processing means and their characteristics in local map construction methods with SDMap.
Following this introduction, the article unfolds as follows: Section II 2 provides a comprehensive review of the pertinent literature, elucidating the general pipeline for local map construction utilizing SDMap. Section III 3 delves into the concept of SDMap, its advantages, and its application scenarios within the task of local map construction. Additionally, we present an overview of the public datasets commonly employed and the evaluation metrics pertinent to tasks of local map perception. Section IV 4 scrutinizes the representation, encoding, and fusion methods of multi-source information, which are crucial for the generation of real-time local maps. This paper further discusses the underlying principles, architecture, and performance of these methods. Section V 5 identifies and discusses key challenges and potential research directions associated with leveraging SDMap as prior information for local map construction. Finally, Section VI 6 synthesizes our findings and offers perspectives on future research endeavors.
@@
Figure 2: General pipeline of local map construction with SDMap, which contains the processing and fusion flow of the three modal data.
2 Related Work
Sensor noise and data delays further complicate the perception task. Therefore, developing robust local map perception technologies is crucial for achieving safe and reliable intelligent driving. To solve this kind of problem, many researchers have proposed various methods. Chen and et al.[1] proposed a method for visual localization and map construction using ground texture, enhancing the accuracy of positioning and the precision of map updates through global and local optimization. SMERF algorithm[2] enhances online map prediction and lane topology understanding by leveraging SDMap and integrating SDMap information through a transformer encoder, which alleviates problems with obscured lane lines or poor visibility and significantly improves the performance of lane detection and topology prediction. [3] proposed an innovative video lane detection algorithm that enhances the feature map of the current frame using an occlusion-aware memory-based refinement (OMR) module, leveraging obstacle masks and memory information to improve the detection accuracy and robustness under occlusion. RVLD[4] improved the reliability of lane detection by recursively propagating the state of the current frame to the next frame, utilizing the information from the previous frames. Besides, there are LaneAF[5], LaneATT[6], and StreamMapNet[7] to ease these issues.
@@ 2 Related Work
2.1 The Task Definition
The task of local map perception involves creating an accurate map representing the vehicle’s surrounding environment to support intelligent driving decision-making and planning. This task typically relies on data from various sensors, including cameras, lidar, radar, and GPS. Additionally, incorporating prior information from SDMap enhances the model’s robustness and mitigates the impact of uncertainties from onboard sensors, improving the overall model performance. The core of the local map perception task is real-time sensing and understanding of the vehicle’s surroundings.
@@ 2.2 Standard Definition Map
SDMap, short for Standard Definition Map, is a digital map technology providing basic geographic information and road network structures. It is widely used in everyday navigation and geographic information services, offering convenience to users. SDMap primarily provides the centerline skeleton of roads without detailed lane information, road signs, or other high-precision environmental features. For the task of local map construction, SDMap offers three main advantages. First, SDMap data is easily accessible. It can typically be obtained for free from open geographic data sources such as OpenStreetMap (OSM)[8], making it suitable for large-scale applications. Second, compared to HDMap, the production and maintenance costs of SDMap are significantly lower. Lastly, SDMap has high universality, covering most types of roads, and can provide relevant road information for local map construction tasks. Platforms like OSM and Baidu map can serve as data sources for SDMap. For instance, OSM is a collaborative project created and maintained by global volunteers, providing free, editable, and open-content maps. OSM data includes a wide range of geographic information such as roads, buildings, parks, and rivers, which can be freely accessed.
@@ 2.3 Datasets
In the field of BEV local map construction, commonly used datasets include KITTI[9], nuScenes[10], ApolloScape[11], Argoverse[12], Openlane[13], and Waymo[14] Open Dataset (Table 1).
Table 1: Datasets for BEV local map construction
dataset Year Locations Scenes Train Val Test Views Map Element HDMap
KITTI[9] 2012 Germany 22 - - - 1 - -
Argoverse[12] 2019 USA 113 200k 40k 80k 7 Only lanes
nuScenes[10] 2020 USA/Singapore 1000 700 150 150 6 -
WaymoOpen[14] 2020 USA 1150 798 202 150 5 Only signs -
CurveLanes[15] 2020 China - 100k 20k 30k 1 - -
ONEC-3DLanes[16] 2022 China - 200k 3000 8000 1 - -
OpenLane-V2[17] 2023 USA/Singapore 2000 1400 300 800 6/7
@@
The KITTI[9] dataset, created by the Karlsruhe Institute of Technology and the Toyota Technological Institute, provides stereo camera, lidar, and GPS/IMU data, covering urban, rural, and highway scenes, and is suitable for tasks such as object detection, tracking, and road detection. The KITTI dataset offers 3D bounding box tracklets for dynamic objects within the camera’s field of view. It defines various classes such as ’Car’, ’Van’, ’Truck’, ’Pedestrian’, and provides annotations for each object’s 3D size, translation, rotation, and occlusion levels.
nuScenes[10], released by Motional, includes data from six cameras, five radars, one lidar, IMU, and GPS, suitable for urban traffic scenarios under various weather and lighting conditions. The nuScenes dataset[10] features comprehensive annotations for intelligent driving, including 3D bounding boxes for 23 classes of objects and 8 attributes. Each scene, lasting 20 seconds, is fully annotated with detailed 3D bounding boxes. The annotation process involves sampling keyframes at 2Hz and is performed by expert annotators to ensure accuracy. The dataset also provides continuous annotations for objects throughout each scene if they are covered by lidar or radar points.
ApolloScape[11], released by Baidu, offers high-precision 3D annotated data covering various urban road scenes, suitable for tasks like lane detection and semantic segmentation. The ApolloScape dataset[11] includes per-pixel semantic segmentation annotations with a unique scale and complexity. It features over 140K images with pixel-level semantic masks and is planned to scale up to 1 million images. The dataset is annotated using an interactive and efficient labeling pipeline that integrates 2D and 3D labeling stages, leveraging high-quality 3D point clouds to enhance the annotation process. Each image is tagged with high-accuracy pose information, and the dataset contains fine-grain labeling of lane markings, making it the first open dataset with 3D annotations for street views.
@@
Argoverse[12], released by Argo AI, includes stereo camera, lidar, GPS, and IMU data, providing detailed 3D annotations and lane markings, primarily used for 3D object detection and lane detection. Argoverse[12] provides detailed 3D tracking annotations, interesting vehicle trajectories, and rich semantic maps. The dataset covers 290 km of mapped lanes with geometric and semantic metadata, using map information to enhance the accuracy of 3D object tracking.
@@
The Waymo Open Dataset[14], released by Waymo, covers a variety of weather and traffic conditions, providing high-quality data from lidar and cameras, suitable for tasks such as 3D object detection, tracking, and lane detection. The Waymo[14] provides comprehensive 2D and 3D bounding box annotations for camera images and lidar data, respectively. The dataset consists of 1150 scenes, each 20 seconds long, with synchronized and calibrated data from urban and suburban areas. Annotations include consistent identifiers across frames to support object tracking.
@@
OpenLane-V2[17], also known as OpenLane-Huawei or Road Genome, is a benchmark dataset for next-generation intelligent driving scene road structure perception, jointly open-sourced by Shanghai Artificial Intelligence Laboratory and Huawei Noah’s Ark Laboratory. It is the first dataset to include the topological relationships of road structures in traffic scenes. OpenLane-V2[17] focuses on topology reasoning for traffic scene structure, offering annotations that describe the relationship between traffic elements and lanes. The dataset comprises 2,000 annotated road scenes that detail traffic elements and their correlation to lanes, aiming to advance research in understanding the structure of road scenes.
@@
ONCE-3DLanes[16] dataset, a real-world intelligent driving dataset with lane layout annotation in 3D space, is a new benchmark constructed to stimulate the development of monocular 3D lane detection methods. It is collected in various geographical locations in China, including highways, bridges, tunnels, suburbs, and downtown, with different weather conditions (sunny / rainy) and lighting conditions (day / night). The whole dataset contains 211K images with its corresponding 3D lane annotations in the camera coordinates.
@@
CurveLanes[15] is a new benchmark lane detection dataset with 150K lane images for difficult scenarios such as curves and multi-lanes in traffic lane detection. It is collected in real urban and highway scenarios in multiple cities in China. All images are carefully selected so that most of them image contains at least one curve lane. More difficult scenarios such as S-curves, Y-lanes, night and multi-lanes can be found in this dataset.
@@ 2.4 Common Evaluation Metrics
The review introduces the evaluation metrics for local map construction methods from two aspects: lane extraction and topology reasoning.
@@ 2.4.1 Metrics for Lane Extraction
Where, mAP represents the mean Average Precision, measuring detection accuracy. The TP set contains the average values of five True Positive Metrics: ATE (Average Translation Error), ASE (Average Scale Error), AOE (Average Orientation Error), AVE (Average Velocity Error), and AAE (Average Attribute Error).
@@ 2.4.2 Metrics for Topology Reasoning
OpenLane-V2[17] has two types of topological reasoning, namely driving scene topology and OpenLane topology. For driving scene topology, OpenLane-V2 defined the OpenLane-V2 UniScore (OLUS), which is the average of various metrics covering different aspects of the primary task. Where $DET_{l}$ represents mAP on lane segments, $DET_{t}$ represents mAP on $DET_{l}$ represents mAP on lane segments and $DET_{a}$ represents mAP on areas. Areas namely pedestrian crossings and road boundaries, are regarded as undirected curves, which are closed or intersected with the boundaries of the BEV range. Chamfer distance is utilized to describe the similarity of areas. $TOP_{ll}$ represents mAP on topology among lane segments and $TOP_{lt}$ represents mAP on topology between lane segments and traffic elements.
$\text{OLUS}=\frac{1}{5}\left[\text{DET}_{l}+\text{DET}_{a}+\text{DET}_{t}+f(\text{TOP}_{ll})+f(\text{TOP}_{lt})\right]$ (4)
For OpenLane topology, OpenLane-V2[17] breaks down the task into three subtasks: 3D lane detection, traffic element recognition, and topology reasoning. The overall task performance is described using the OpenLane-V2 Score (OLS), which is the average of the metrics for each subtask. The metric for 3D lane detection, $DET_{l}$, can be expressed as the mean AP at different thresholds $t\in T$, $T$={1.0, 2.0, 3.0}, where AP is calculated using the Fréchet distance. Traffic element detection is evaluated similarly to object detection using AP, with an IoU threshold set to 0.75. Traffic elements have various attributes, such as traffic light colors, which are closely related to lane accessibility, so attributes must also be considered. Assuming A is the set of all attributes, the evaluation includes attribute classification accuracy.
$\text{OLS}=\frac{1}{4}\left[\text{DET}_{l}+\text{DET}_{t}+f(\text{TOP}_{ll})+f(\text{TOP}_{lt})\right]$ (5)
@@
OpenLane-V2[17] uses the TOP score to evaluate the quality of topology reasoning, akin to the mAP metric but adapted for graphs. Essentially, this converts the topology prediction problem into a link prediction problem and calculates mAP (mean APs of all vertices) to assess algorithm performance. The first step is to determine a matching method to pair ground truth and predicted vertices (i.e., centerlines and traffic elements). For centerlines, Fréchet distance is used; for traffic elements, IoU is used. When the confidence score of an edge between two vertices exceeds 0.5, they are considered connected. The vertex AP is obtained by ranking all predicted edges of a vertex and calculating the mean of cumulative precision: